jhonyy9 (jhonyy9):

true or false ? sec x ------- cosec x -------- = 1 tan x

jhonyy9 (jhonyy9):

bc. we know that sec = 1/cos and cosec = 1/sin so sec/cosec = 1/cos / 1/sin = (1/cos) *(sin/1) = sin /cos = tan so tan / tan = 1

OpenStudy (moldybubblegum11):

i think its true

Nnesha (nnesha):

good job!

OpenStudy (wckeller):

Correct.

jhonyy9 (jhonyy9):

ty Nnesha

Nnesha (nnesha):

yw o^_^o